Laura’s Notebook: Retention Time R2 Calcs

One of the first steps in processing SRM data is to confirm that the selected peaks actually represent the peptides, aka that our assay works. To do this, we use linear regression between PRTC retention times in DIA and SRM to calculate predicted transition RTs in collected SRM data. Then, we calculate the R^2 for PRTC and experimental peptides compared to predicted.

Step 1

Copied PRTC peptides from another Skyline project file, pasted into my Geoduck assay Skyline file. image

Step 2

The method file used in mass spec run resulted in us not collecting all PRTC peptides, so I removed those from the Skyline document. snip20170811_2

Some of the PRTC peptides had no signal in several of my replicates. These peptides eluted at ~14 or 18 minutes. I removed the peaks from all reps which had no signal at the predicted RT. In this screen shot the peptides with lots of signal missing are unfolded on the left side, and one of the cruddy reps’ chromatogram is pulled up. There should be a peak @ around 17mins. image

Notice that the PRTC peptide signals vary, which is due to me using different PRTC mixes. I documented which mix I used, and the concentrations of each, so I will need to account for that when I normalize my assay daya using PRTC data. More on that later.

Step 3

Export retention times from Skyline via File -> Export -> Report -> Peptide RT Results. Saved in my Geoduck-DNR Repo as 2017-08-11-SRM-Retention-Times.csv image

more in a minute…

from LabNotebook http://ift.tt/2uyU4lL
via IFTTT

Laura’s Notebook: Cleaning up SRM data in Skyline, part II

Before I export my data from Skyline for data analysis, I have the following final things to do:

  • Adjust peak boundaries
  • Determine if any .raw files should be discarded, based on poor-quality data and those that I re-ran & re-made
  • Look @ all blank runs to see if there are any weird signals
  • Check out blank samples

Adjusting peak boundaries

Some helpful keyboard shortcuts:

  • Scroll between replicates: Ctrl+Up or Ctrl+Down
  • Auto-zoom to best peak: F11
  • Un-autozoom out from best beak: Shift+F11

Some transition peaks look poor, but they are present. Here is an example of replicate data for the Superoxide Dismutase protein, showing the overall view and the zoomed view: imageimage

This is in comparison to the following replicate, where there is no peak present betwen RT 14-15: imageimage

Not sure what to do in the situation where a peak is split into 2 peaks (as below); Skyline opted for the boundaries to encompass both peaks. I will do the same, as the total RT for both peaks appears to be similar to that of other reps: image

Notes:

  • Poor quality reps: 178, 254, 208, 212, 213, 297_170728020436,
  • A peptide with RT ~18 must be co-eluting, as it pops up in a few res/peptides. Perhaps it is the [PEPTIDE W/ 18], that gets stuck in the column and pops up. For example, the following is a zoomed-out view of Ras-related protein peptide, which should have it’s peak around 22.7. This is rep #254, but it pops up in lots of reps: imageimage A couple peptides elute @ ~18min, and could be the culprit: Sodium/potassium-transporting ATPase subunit alpha-4, MVTGDNVNTAR; Catalase, LYSYSDTHR;

Total actions performed on SRM data in Skyline:

  • Removed peaks from replicates where no peak was present @ designated retention time (as per DIA/SRM regression). Replications w/ no peaks for peptides will be represented as #N/A when data is exported.
  • ID’d peptides with very poor data across multiple replicates; I may not use these peptides in my analysis; TBD. See peptides in red
  • Adjusted retention time boundaries for all reps all peptides. I erred on NOT adjusting boundaries if they looked OK to maintain consistency.
  • ID’d and deleted 2 transitions that do not align with other transitions at designated RT. Transitions are:
    • Superoxide dismutase, TIVVHADVDDLGK, y4
    • Ras-related protein Rab-11B VVLVGDSGVGK, y4
  • Documented peptides and transitions w/ poor quality over several samples here

Exported data from Skyline:

Export -> Report, then I edited the Transition Results report with the following metrics: Protein Name, Transitions, Peptide Sequence, Fragment Ion, Peptide Retention Time, Area; I selected “Pivot Replicate Name”. Here’s a preview of the report: image

from LabNotebook http://ift.tt/2vv8Dbg
via IFTTT

Laura’s Notebook: My oysters are growing up

Screened all oysters up to 450um screen, counted set on 180um silos

I set my oysters on 224um microcultch, and hold them on 180um silos until they grow enough to hold on 450um, which should happen within a week or so. Until today I had a dual upwell/downwell system:

upwell/downwell systemupwell/downwell system2

The last 224um larvae went into 180um setting tanks on 7/20, 3 weeks ago. Successful set should therefore all be at least 450um. Today I screened all 180um silos on to 450um, inspected under the microscope for seed, and moved them to the upwelled silos. I inspected a few of the <450 contents under the scope and found zero live seed. I also moved all SN growth experiment seed from mini-silos into their own 450um upwelled silo. Lots of oysters set on the sides of the silos. I counted these today. I also was able to pull a few off the silos and added them to the upwellers. I wasn’t able to get through all the silos, so I’ll do so on Saturday and will provide total # salvaged then. seed set on silo

Also counted the # set on side of SN growth exp silos (none salvaged): – 6-Ambient = 3 – 6-Low = 18 – 10-Ambient = 0 – 10-Low = 1

from LabNotebook http://ift.tt/2vTV875
via IFTTT

Yaamini’s Notebook: Gigas Larvae Day 9

Totes are pesky

I walked into the SeaLab this morning and heard the sound of an alarm. I saw that Tote 7 was reading at 21 ºC, which was quite low. When I looked into the tote, I saw that it was completely empty, and the plastic casing around the heater had melted!

img_8222

img_8223

Figures 1-2. Empty Tote 7 and melted heater.

I filled the tote with heated water to track down where the water had leaked from, and to keep the buckets a little warmer while I looked for a new tote. I found another smaller black tote in the warehouse, similar to the two blacks ones I was already using. I cleaned it and filled it with heated seawater. Because it could only hold two buckets, I moved Bucket 10 to Tote 1. Luckily, Tote 7 had an AVTECH in it and Buckets 3, 10 and 19 had HOBOs. I was confused because I didn’t get any alert about the temperature in this tote dropping. When I checked the Dashboard, I realized I did not set alarms for Gigas-Larvae-3 or Gigas-Larvae-4. I set those alarms up in case something like this happens again.

AVTECH inspection

I looked at the AVTECH and saw that around 10 p.m. the water temperature started dropping, reaching 20ºC in the tote this morning.

screen shot 2017-08-08 at 10 49 13 am

Figure 3. AVTECH Dashboard reading of temperatures.

While looking at the Dashboard, I noticed that Gigas-Larvae-2 dropped in temperature early this morning and then started climbing.

screen shot 2017-08-08 at 10 59 09 am

Figure 4. AVTECH Dashboard reading.

This AVTECH is in Tote 1, which had an additional bucket added to it. The temperature in the Bucket 10 was lower than the tote temperature, causing the decrease in temperature. The heater is now working to bump the temperature back up.

I also noticed with the AVTECHs is that Gigas-Larvae-3 and Gigas-Larvae-4 had stopped reporting temperatures on the Dashboard after 8 a.m. this morning. They both come from WISH Ext. 9. I checked the computer in the dry lab and saw that both AVTECHs were reporting at 24ºC on that monitor, so maybe there’s just a problem with that information being pushed or synced properly to the Dashboard.

Finally, I found another AVTECH tucked behind one of the totes! I identified it as Temp 2 and changed the name on the dashboard to Gigas-Larvae-Bucket. I thought it would be a good idea to see how the temperature differs between the tote and the buckets inside, so I placed the AVTECH in Tote 5, Bucket 12. Based on these readings, I can adjust the temperature of the immersion heaters if needed. I didn’t set any alerts since I maxed out the number I could have, but at least I can monitor what happens.

Power outage

The hatchery power went out at 11:30 a.m. My immersion heaters were on non-emergency power so they were working fine. However, my airline stopped working and the AVTECH Wifi also went out. At 4 p.m., power was restored. The larvae went 4.5 hours without any bubbling. I adjusted the bubbling of the airlines and checked all temperatures on the AVTECH. Temperatures looked good around 24ºC!

Other tasks

I talked to Stuart, who said he could seal the crack in the that tote and the other tote that cracked before even starting the experiment with marine epoxy. These will be good to have as a back-up, but I will also order some more totes just in case. The totes I have are Rubbermaid brand and can hold 54 gallons. I want totes that are a bit thicker on the bottom and can fit 3-5 gallon buckets. I will also order a few more immersion heaters since this one isn’t useful anymore.

After moving things around, I recorded tote positions for all buckets. I have an entry for tote positions after yesterday’s water change, and then the slightly modified tote positions from this morning.

I added HOBOs to Buckets 4, 6, 7, 8, 14, 16, 18, 20, 22 and 24, since I didn’t have HOBOs in them before. I checked that they had enough battery life, then set them to record data every 30 minutes. I may change this tomorrow. When I added the HOBOs to the buckets, I noticed Bucket 17 and 23 no longer had HOBOs in them, and the HOBO from Bucket 2 was sitting on the table! I added Bucket 2’s HOBO back in at 12:43 p.m. I put two new loggers in 17 and 23 and searched for the original ones, but couldn’t find them. Maybe someone put it on the side somewhere while screening? I’ll figure it out!

While waiting for the power to turn back on, I went over to ACE Hardware in Port Orchard. They didn’t have any tubs big enough to hold more than one five-gallon bucket. I found these totes on Amazon that will work, and can be delivered in two days. I created a new purchasing issue for the totes and heaters.

I fed my larvae around 1 p.m. so they weren’t completey without food when the power was out. I prepped tripours for tomorrow’s screening, but I was almost out of sample tubes. I’ll need to grab more from the lab before tomorrow morning.

Just as I was about to leave, Tote 9’s high alarm went off. As I was emptying the water and refilling it, the heater’s probe read at 30ºC. I don’t know what it is about that tote or immersion heater that always sets the alarm off. I was able to get the temperature down to 26ºC in the tote after 2-3 minutes.

Help schedule for this weekend

  • Thursday: Kelsey feeding
  • Friday: Grace, Kelsey and Ashley screening, counting, feeding, sampling, draining and refilling header tank
  • Saturday: Laura feeding
  • Sunday: Grace and Kaitlyn screening, counting, feeding, sampling, draining and refilling header tank

I’m sure I can make it out to Manchester on Monday, but just in case my flight is delayed, I will ask someone to feed.

For tomorrow

  • GET SAMPLE TUBES FOR SAMPLING LARVAE
  • Screen on 80 micron and 60 micron screens
    • Measure amount to feed ahead of time
    • Add food while restocking larvae
    • Top totes off with freshwater, not heated seawater
  • Drain header and refill
  • Prepare tasklists for Grace (Friday and Sunday)

from yaaminiv.github.io http://ift.tt/2vBPoPW
via IFTTT

Laura’s Notebook: Cleaning Srm Data In Skyline

First steps, done on 8/8

Clean up data in Skyline:

Used final version of sequence file (which I uploaded from the Vantage computer to Owl @ UWPR on 7/28) to associate .raw files with samples – determine which samples are which for the duplicate named files. The final sequence file, which will also contain the time loaded into tray & injection time (in case this is a factor), is located in the Data folder of this repo, here.
image

Make sure I have all data in Skyline

Confirmed in Skyline by going to Edit -> Manage Results & cross-checking data file names in Skyline with the sequence file, and the .raw files on Owl. image

Confirm correct peaks picked & determine if any transitions/peptides are un-usable

Referenced the Predicted SRM Retention Time calculations I did at the beginning of my mass spec run to confirm correct peaks picked. ID’d peptides with lots of noice, and no clear peak at the predicted RT. Here is a screen shot of the predicted RT, with poor quality peptides in red. NOTE: Need at least 2 transitions for each peptide to move forward with analysis. image

Next steps:

  • Adjust peak boundaries
  • Determine if any .raw files should be discarded, based on poor-quality data and those that I re-ran & re-made
  • Look @ all blank runs to see if there are any weird signals
  • Check out blank samples
  • Download data! Protein, peptide, transition, area, retention time
  • Work through Emma’s suggestions: normalization, dilution durve check, NMDS, ROC curves, data upload

from LabNotebook http://ift.tt/2uEDZtY
via IFTTT

Yaamini’s Notebook: Gigas Larvae Day 8

Immersion heaters are pesky

Overall, today was pretty smooth. Olivia and Grace helped me screen my larvae. At first, I thought I would be able to screen on both an 80 micron and 60 micron screen. However, after screening the first two buckets, Bucket 11 and 6, I didn’t count any larvae from the 80 micron screen. I decided to continue just screening on a 60 micron screen, and try again with two screens on Wednesday. I did something different today where I added food to the bucket before restocking, instead of adding food to all buckets at the end. I think this method is better because the larvae get fresh algae after being stressed by the screening process, so I’ll continue to do this. While screening Bucket 12, Grace accidentally screened through a 48 micron screen first. She caught her mistake very quickly, but it’s possible some larvae were lost with this additional screen. All of my larvae are at about 2 larvae/mL, which is the stocking density I want at this point. You can see my counts and feeding log here.

Now here’s the annoying part. Totes 3, 4, 5, 7 and 9 all had their high alarms (27ºC) go off after the buckets were restocked. Tote 4 and 7 had their alarms go off multiple times! To reduce temperature, we used tripours to drain out some of the water in the tote, and then we refilled with ambient freshwater instead of using the water from the heating reservoir line. This quickly brought our temperatures down. I think refilling totes with ambient freshwater should be our standard now. Adding heated water to a tote that’s already at temperature may set it over the edge.

For tomorrow:

  • Record tote positions
  • Add new HOBO loggers buckets without them
  • Feed oysters
  • Finalize help schedule for this weekend
  • Prepare task lists for those helping

from yaaminiv.github.io http://ift.tt/2vjjIMD
via IFTTT

Yaamini’s Notebook: Gigas Larvae Day 7

Anoxia is no bueno

My main goal today was to flush out any anoxic water from the heated line. When I got to the hatchery, I started draining the header tank and opened the heated line to flush any water out. I let this run while I prepared tripours for tomorrow’s screening and fed Joth’s scallops. After 30 minutes, I shut the heated line off and replaced the 1 micron filter. I noticed that the filter casing didn’t have the black rubber ring around it anymore, but water wasn’t leaking from it. I rinsed the casing and put a new filter in, then turned the heated line back on. I let the line run for 2 minutes while I took a freshwater hose and cleaned the inside of the header tank. After a few more minutes of letting the tank drain, I closed the bottom valve and started filling it.

While the tank was filling, I fed my larvae. I noticed that the buckes were relatively clear, and most of the algae had settled to the bottom.

img_8203-1

Figure 1. Tank color upon arrival.

I figured I would increase the algal density to about 65,000 cells per mL and check the color of the buckets tomorrow morning before screening. I don’t want the larvae to go without food dispersed throughout the water for more than an hour, so maybe increasing the amount of food will help maintain the algae level I want. It’s also possible they’ve started to eat more as they’re growing.

After feeding my larvae, the header tank was about halfway filled. I plugged all three heaters back in, kept the heated line open so the tank could finish filling, and rushed out the door to catch the 11 a.m. ferry back. In my haste, there were a few small screening preparation tasks I forgot to do, so they’re at the top of my list tomorrow!

For tomorrow:

  • Record which buckets are in which totes (I need this before I rerandomize my buckets!!)
  • Turn off cold water input
  • Label tubes for larval samples
  • Measure out amount to feed ahead of screening, so larvae can be fed while being restocked
  • Screen larvae on 80 and 60 micron screens
  • Count larvae and drop stocking density to approximately 2 larvae/mL
    • Based on my current count data, Buckets 3, 9, 10 and 11 are the four that are above this stocking density
  • Drain heating header and unplug heaters
  • Clean and/or replace filters
  • Return Rick’s silver heater

from yaaminiv.github.io http://ift.tt/2fjPJQo
via IFTTT